OpenStudy (anonymous):

The midpoint of a segment is (2,-5) and one of the endpoints is (3,6) where is the other endpoint?

OpenStudy (anonymous):

DId you get an answer yet?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Well since (2, -5) is your midpoint and one of your endpoints is (3, -6), where do you think you the other endpoint is or how do you think you can get it.

OpenStudy (anonymous):

whoops my bad. Your endpoint is (3, 6). So can you think of a way that might work?

OpenStudy (mathstudent55):

For points (x1, y1) and (x2, y2), the midpoint is ( (x1 + x2)/2, (y1 + y2)/2 ) You have (x1, y1) = (3, 6) and midpoint (2, -5), so (3 + x2)/2 = 3 (6 + y2)/2 = -5 Solve the above equations for x2 and y2

OpenStudy (anonymous):

Well mathstudent your equation is wrong for the first one. It should = 2